# Adapting Vision-Language Models for Evaluating World Models

### v1

@@ Abstract
World models – generative models that simulate environment dynamics conditioned on past observations and actions – are gaining prominence in planning, simulation, and embodied AI. However, evaluating their rollouts remains a fundamental challenge, requiring fine-grained, temporally grounded assessment of action alignment and semantic consistency – capabilities not captured by existing metrics. Vision-Language Models (VLMs) have shown promise as automatic evaluators of generative content due to their strong multimodal reasoning abilities. Yet, their use in fine-grained, temporally sensitive evaluation tasks remains limited and requires targeted adaptation. We introduce a evaluation protocol targeting two recognition tasks – action recognition and character recognition – each assessed across binary, multiple-choice, and open-ended formats. To support this, we present universe (UNIfied Vision-language Evaluator for Rollouts in Simulated Environments), a method for adapting VLMs to rollout evaluation under data and compute constraints. We conduct a large-scale study comparing full, partial, and parameter-efficient finetuning across task formats, context lengths, sampling strategies, and data compositions. The resulting unified evaluator matches the performance of task-specific baselines using a single checkpoint. Human studies confirm strong alignment with human judgments, establishing universe as a scalable, semantics-aware evaluator for world models.
@@ 1 Introduction
World models are generative models trained to predict future observations conditioned on past observations and actions [37, 41, 5]. They offer a powerful abstraction for learning, reasoning, and planning in complex interactive environments, and are rapidly becoming foundational in domains such as neural game engines [56, 35, 33, 21], embodied AI [31, 106], and autonomous driving [84, 47, 77]. As their capabilities grow, a persistent challenge remains: evaluation.
Rollouts produced by world models are visually rich, temporally grounded, and semantically structured. Evaluating them requires: (i) precise alignment between generated frames and control sequences at the timestamp level [106], and (ii) consistent tracking of entities and their semantics over time [56].
Existing evaluation protocols offer limited insight in this regime. Traditional visual generation metrics are ill-suited to the structured and dynamic nature of rollouts: (i) early distributional metrics focus on static images and are sensitive to low-level variations [85, 46, 12], (ii) motion-aware metrics such as FVD [97] capture temporal patterns but lack semantic grounding, and (iii) multimodal metrics improve semantic alignment but ignore action conditioning and timestamp-level fidelity [52]. While human evaluation remains the gold standard [2, 6], it is expensive and difficult to scale. Similarly, emerging text-to-video benchmarks [66, 50, 62] focus on open-ended generation and neglect the fine-grained control and temporal coherence essential to evaluating world model outputs.
[…]
VLM have demonstrated strong generalization across multimodal tasks by integrating visual and linguistic reasoning [61, 30, 24, 99, 1, 64, 26, 72], and are increasingly used to evaluate generative models [59, 70, 63, 20]. We extend this direction, hypothesizing that VLMs can serve as fine-grained evaluators of world model rollouts, capturing semantic consistency, temporal coherence, and alignment with control inputs. Virtual environments offer a favorable testbed for this purpose, as they expose full access to timestamped actions, object states, and agent behaviors, enabling controlled and systematic evaluation.
@@
Figure 1: Performance and efficiency of universe (orange bars throughout) compared to task-specific baselines (multiple colours). Left and Center: Action recognition and Character Recognition accuracy across binary, multiple-choice, and open-ended settings. Right: Sample efficiency – our adaptation recipe achieves strong performance with substantially fewer training samples per epoch. universe matches the performance of task-specific baselines while using a single checkpoint, enabled by mixed supervision, efficient frame sampling, and lightweight partial fine-tuning.
[…]
To support this protocol, we introduce universe (UNIfied Vision-language Evaluator for Rollouts in Simulated Environments), a method for adapting VLMs to structured rollout evaluation. universe emerges from a systematic study of adaptation strategies under realistic data and compute constraints. We analyze the impact of supervision regime, frame sampling strategy, visual context length, and training budget, arriving at an adaptation recipe that combines mixed supervision, efficient frame selection, and lightweight fine-tuning.
To assess the reliability of universe, we conduct a human annotation study using rollouts generated by WHAM [56], a publicly available world model trained in a complex interactive environment. We evaluate the rollouts using universe and collect human judgments on the same data to compare responses. We find that universe’s judgments exhibit strong alignment with human ratings, confirming its potential for automated evaluation world model rollouts, particularly when explicit ground truth is unavailable or costly to obtain.
In summary, this work makes three key contributions:
  1. I
    We introduce a structured protocol for evaluating world model rollouts, based on action and character recognition tasks with increasing levels of complexity;
  2. II
[…]
  3. III
    We show that universe matches the performance of task-specific baselines using a single checkpoint, demonstrating its ability to generalize across tasks with a unified model (Figure 1). Furthermore, its predictions align closely with human judgments, validating its effectiveness as a scalable, semantics-aware evaluator for world model rollouts.
@@ 2 Related Work
Challenges in Evaluating World Models. World models are generative systems that learn predictive representations of environment dynamics [37], originally proposed for model-based RL [92] and now central to domains such as neural game engines [56, 35, 33, 21], embodied AI [31, 106], and autonomous driving [84, 47, 77]. Recent models such as Dreamer v1–3 [39, 38, 40, 41], MuZero [86], IRIS [74], UniSim [106], and DIAMOND [5] have improved rollout fidelity and controllability. Yet evaluation largely focuses on downstream success metrics—e.g., game score or goal completion [9, 54, 36, 7, 8]—which provide only coarse, indirect signals of rollout quality. Genie [17, 78] decouples world model learning and agent training, but its evaluation still emphasizes visual quality and control, without probing semantic or causal fidelity. Cosmos [2] proposes a structured protocol that combines FID/FVD with structure-from-motion-based 3D consistency checks and human ratings on instruction following, object permanence, and visual verity. While insightful, this approach is tied to simulator-specific infrastructure and requires costly manual comparison. Human-in-the-loop protocols such as the Video Generation Arena [6] also rely on pairwise comparison to assess rollout quality. These methods, though informative, are expensive and hard to scale.
@@
Evaluation Metrics and Protocols for Visual Generation. Early evaluations of generative models relied on full-reference metrics such as PSNR and SSIM [100], which capture pixel-level and perceptual similarity but are sensitive to spatial misalignments and fail to reflect semantic fidelity. To address this, distributional metrics like Inception Score (IS) [85], Fréchet Inception Distance (FID) [46], and Kernel Inception Distance (KID) [12]. Other proposals such as PPL [57], Parzen likelihoods [34], and HYPE [110] attempt to quantify perceptual smoothness or human realism, but remain focused on static images. For video generation, FVD [97] generalizes FID using I3D features [18], introducing a motion-aware distributional baseline. Yet, FVD also lacks semantic grounding and does not account for causal structure or goal alignment. To improve semantic grounding, metrics based on text-image alignment have been proposed. CLIPScore [45] and CLIPSIM [102] compute similarity between generated visuals and textual or visual references using CLIP embeddings [82], while Jayasumana et al. [52] extend this to distributional comparisons via MMD. However, all operate at the frame level. Structured evaluation protocols using vision-language reasoning have also emerged. VQA Accuracy [70] uses LLMs to score answers on static image questions, and VQAScore [63] probes alignment via templated binary queries. Lee et al. [59] propose VLM evaluator to evaluate other VLMs responses given user criteria. These approaches introduce task structure but remain limited to single-frame evaluation. Recent text-to-video (T2V) benchmarks such as EvalCrafter [66], VBench [50], and DEVIL [62] introduce curated prompts and metrics covering text alignment, motion realism, and perceptual quality. While these protocols push forward evaluation of open-ended video generation, they lack timestamp-level action grounding.
@@
Vision-Language Model Adaptation. VLMs have emerged as powerful tools for multimodal understanding, demonstrating strong performance across tasks such as captioning, retrieval, visual question answering, and instruction following [61, 30, 24, 99, 1, 64, 26, 72]. Adaptation approaches can be broadly categorized into prompt-level and weight-level methods. One prominent prompt-level adaptation techniques is prompt tuning, which injects task information directly into the input space [75, 111, 103], and in-context learning (ICL), where models such as GPT-3 [16] and Flamingo [4] condition on task demonstrations at inference time without updating parameters. Retrieval-augmented generation (RAG) [60] combines parametric models with non-parametric memory, and multimodal variants incorporate external visual or auditory context [49, 22]. While lightweight, these approaches are limited in their ability to model temporal dependencies or align with structured rollouts. Weight-level adaptation enables stronger domain alignment but incurs higher computational cost. Full finetuning remains effective yet costly, while partial finetuning [107] offers a trade-off by updating only selected layers. Parameter-efficient finetuning (PEFT) provides a scalable alternative and can be grouped into low-rank and adapter-based strategies [42]. Low-rank methods, such as LoRA [48], inject rank-constrained updates into frozen layers. Recent extensions improve upon this via weight decomposition [65], quantization-aware adaptation [27, 105], mixture-of-experts routing [104], and long-context support [25]. Adapter-based methods insert lightweight modules between frozen layers to enable modular adaptation with minimal overhead [69, 109]. A parallel line of work investigates multimodal few-shot learning. Frozen [95] was among the first to explore this setting, followed by works combining prompting and ICL for improved sample efficiency [53, 89], and works introducing a learnable meta-mapper to bridge frozen VLM components for few-shot meta-learning [76].
[…]
@@ 3 Methodology
[…]
Dataset Construction. Effective VLM adaptation for rollout evaluation requires a dataset that (i) captures realistic human behavior in interactive environments, and (ii) aligns with prior work in simulated settings to support comparability and reproducibility. Additionally, since WHAM – the world model used in our evaluation – is trained on Bleeding Edge, it is critical that the adaptation dataset match its distribution. To satisfy these constraints, we partnered with Ninja Theory and curated a dataset from both internal and public Bleeding Edge gameplay recordings, focusing on the Skygarden map used in WHAM [56]. This dataset provides high visual and behavioral diversity [79], includes a publicly available evaluation split, and is closely aligned with prior work in the domain [56, 79, 94, 88, 28], enabling cross-method comparison.
Data preparation proceeds in three stages: (i) Preprocessing:Segment gameplay into 14-frame clips with synchronized video, control logs, and metadata; (ii) Description Generation:Convert structured annotations (e.g., actions, agent states) into natural language summaries; (iii) Question-Answer Pair Construction:Generate six QA pairs per clip (binary, multiple-choice, and open-ended) spanning both AR and CR tasks. The final dataset contains 32.453 training clips and 8.113 validation clips, yielding 194.718 and 48.678 QA pairs, respectively. See Appendix C for details.
Model Architecture. We adapt a model from the PaliGemma family [10, 91], consisting of a vision encoder $\mathcal{M}_{V}$, a projection head $\mathcal{M}_{P}$, and a language decoder $\mathcal{M}_{L}$. Based on initial zero-shot evaluations (Appendix F.1), we use a single configuration for all experiments—PaliGemma 2 3b, which includes a 2B-parameter Gemma 2 decoder pretrained on 2T tokens. Input frames are resized to $224\times 224$ and tokenized into 256 patches each. Model achitecture details are in Appendix D.1.
Each model input sequence $S=\{S_{\mathcal{I}},S_{\mathcal{T}}^{\text{PREF}},S_{\mathcal{T}}^{\text{SUFF}}\}$ consists of: visual tokens $S_{\mathcal{I}}$ from $k$ frames, a textual prefix $S_{\mathcal{T}}^{\text{PREF}}$ containing the task-language cue and question, and a suffix $S_{\mathcal{T}}^{\text{SUFF}}$ with the expected answer (used only during training). This format allows the decoder to attend jointly over visual and textual context. Full prompt details are provided in Appendix D.1.
@@
Fine-Tuning Configurations. We compare five adaptation strategies varying in parameter count and modularity: (i) Zero-shot prompting:No tuning; model is prompted directly. (ii) Full fine-tuning:All parameters $\theta=\theta_{V}\cup\theta_{P}\cup\theta_{L}$ are updated end-to-end. (iii) Dual-component fine-tuning:Two of three modules are trained (e.g., $\theta_{P}\cup\theta_{L}$). (iv) Single-component fine-tuning:Only one module—vision, projection, or language—is updated. (v) Parameter-efficient fine-tuning:We apply LoRA [48] adapters to attention and MLP layers in vision and language components: $\mathbf{W}\leftarrow\mathbf{W}+\frac{\alpha}{r}\mathbf{A}\mathbf{B},\quad\alpha=8,\;r\in\{8,16,32,48,64\}$.
@@
universe: UNIfied Vision-language Evaluator for Rollouts in Simulated Environments. We distill our empirical findings into universe, a lightweight and scalable adaptation method for temporally grounded evaluation of world model rollouts using VLMs. Designed for constrained compute and limited supervision, universe delivers strong generalization across our evaluation protocol using a single, partially tuned model. The method combines three key components:
  1. I
    Partial fine-tuning: We update only the projection head ($\theta_{P}$), training just 0.07% of model parameters. Despite this minimal footprint, it achieves the second-best performance among all strategies—trailing only vision encoder tuning, which requires $\sim$11of parameters and incurs significantly higher compute cost.
  2. II
    Efficient frame sampling: Each input sequence includes $k=8$ frames sampled uniformly from a 14-frame rollout. This sparsity-aware strategy maintains long-range temporal structure while reducing token count and enabling efficient batching.
  3. III
    Mixed supervision: We train on a hierarchical mixture of tasks and QA formats. The task distribution favors Action Recognition ($\alpha_{\texttt{AR}}=0.8$) due to its stronger causal grounding. Within each task, we emphasize open-ended questions ($\beta_{\texttt{OE}}=0.8$), while maintaining smaller proportions of binary ($\beta_{\texttt{binary}}=0.15$) and multiple-choice ($\beta_{\texttt{MC}}=0.05$) examples.

### v2

@@ Abstract
World models – generative models that simulate environment dynamics conditioned on past observations and actions – are gaining prominence in planning, simulation, and embodied AI. However, evaluating their rollouts remains a fundamental challenge, requiring fine-grained, temporally grounded assessment of action alignment and semantic consistency – capabilities not captured by existing metrics. Vision-Language Models (VLMs) have shown promise as automatic evaluators of generative content due to their strong multimodal reasoning abilities. Yet, their use in fine-grained, temporally sensitive evaluation tasks remains limited and requires targeted adaptation. We introduce an evaluation protocol targeting two recognition tasks – action recognition and character recognition – each assessed across binary, multiple-choice, and open-ended formats. To support this, we present universe (UNIfied Vision-language Evaluator for Rollouts in Simulated Environments), a VLM-based evaluator for video world model rollouts adapted under data and compute constraints. In our extensive experiments totaling over 5,154 GPU-days, we explore full, partial, and parameter-efficient adaptation methods across various task formats, context lengths, sampling methods, and data compositions. The resulting unified evaluator achieves parity with task-specific checkpoints. Human studies across seven diverse environments confirm strong alignment with human judgments, establishing universe as a lightweight, adaptable, and semantics-aware evaluator for video world models.
@@ 1 Introduction
World models – generative models trained to predict future observations conditioned on past observations and actions (Ha & Schmidhuber, 2018; Hafner et al., 2025; Alonso et al., 2024) – are rapidly becoming central to interactive AI. They provide a powerful abstraction for learning, reasoning, and planning in complex interactive environments, and underpin advances in neural game engines (Kanervisto et al., 2025; Guo et al., 2025; Gao et al., 2025; Chen et al., 2025), embodied AI (Du et al., ; Yang et al., 2024), and autonomous driving (Russell et al., 2025; Hu et al., 2023a; Ni et al., 2025).
Yet, evaluating world models remains a bottleneck. Rollouts are semantically rich and temporally grounded, requiring metrics that assess (i) alignment between generated frames and action sequences at the timestamp level (Yang et al., 2024), and (ii) consistent entity tracking over time (Kanervisto et al., 2025). Existing approaches fall short: (i) early distributional metrics focus on images and are sensitive to low-level variations (Salimans et al., 2016; Heusel et al., 2017; Binkowski et al., 2018), (ii) motion-aware metrics like FVD (Unterthiner et al., 2018) lack semantic grounding, and (iii) multimodal metrics ignore timestamp-level action conditioning (Jayasumana et al., 2024). Emerging text-to-video benchmarks (Liu et al., 2024b; Huang et al., 2024; Liao et al., 2024) focus on open-ended generation but neglect the fine-grained control central to world model evaluation. Even cutting-edge LLMs fail in this setting (Appendix G.1, Figure 15). Human evaluation remains the gold standard (Agarwal et al., 2025; Analysis, 2024), however, it remains costly and hard to scale.
To address this gap, we propose a novel evaluation protocol targeting two important dimensions of rollout quality: action alignment and character consistency, formalized as recognition tasks – Action Recognition (AR), Character Recognition (CR) – across formats of varying complexity. The protocol provides a foundation for semantic evaluation of rollouts, and extends upon existing evaluation method by providing insight into semantic quality of generated rollouts. Inspired by the LLM-as-a-Judge research direction Zheng et al. (2023), we explore whether Vision-Language Models can serve as effective evaluators in this setting. VLMs have demonstrated strong generalization across multimodal tasks (Li et al., 2023; Driess et al., 2023; Chen et al., 2023; Wang et al., 2024; Abdin et al., 2024; Liu et al., 2024a; Deitke et al., 2024; McKinzie et al., 2024), and show promise as automatic judges of generative models (Lee et al., 2024; Mañas et al., 2024; Lin et al., 2024; Chen et al., 2024). Yet, off-the-shelf VLMs struggle with the temporal grounding and domain-specific knowledge (see Sec. 4, Zero-Shot Evaluation). Limited resources and sparse text supervision introduce another layer of complexity to the setting.
We therefore conduct a systematic study of adaptation strategies under realistic data and compute constraints. Throughout the study totaling over 5,154 GPU-days, we analyze the impact of supervision regime, frame sampling strategy, visual context length, and training budget. The result of the study is universe (UNIfied Vision-language Evaluator for Rollouts in Simulated Environments), a lightweight, adaptable, and semantics-aware evaluator for world models. universe achieves parity with six task-specific checkpoints while using a single unified model (see Figure 1. To validate reliability, we conduct a large-scale human study on rollouts spanning diverse environments, model scales, and rollout fidelities. universe’s judgments show strong agreement with human ratings, demonstrating its effectiveness as a practical, semantics-aware evaluator—particularly valuable in settings where ground-truth annotations are unavailable or prohibitively expensive.
@@
Figure 1: Performance and efficiency of universe (orange bars throughout) compared to task-specific baselines (multiple colours), all models trained for 10 epochs. Left and Center: Action recognition and Character Recognition accuracy across binary, multiple-choice, and open-ended settings. Right: Sample efficiency – our adaptation recipe achieves strong performance with substantially fewer training samples per epoch.
@@ 2 Related Work
Challenges in Evaluating World Models. World models are generative systems that learn predictive representations of environment dynamics (Ha & Schmidhuber, 2018), originally proposed for model-based RL (Sutton, 1991) and now central to domains such as neural game engines (Kanervisto et al., 2025; Guo et al., 2025; Gao et al., 2025; Chen et al., 2025), embodied AI (Du et al., ; Yang et al., 2024), and autonomous driving (Russell et al., 2025; Hu et al., 2023a; Ni et al., 2025). Recent models such as Dreamer v1–3 (Hafner et al., 2020; ; 2023; 2025), MuZero (Schrittwieser et al., 2020), IRIS (Micheli et al., 2023), UniSim (Yang et al., 2024), and DIAMOND (Alonso et al., 2024) have improved rollout fidelity and controllability. Yet evaluation largely focuses on downstream success metrics—e.g., game score or goal completion (Bellemare et al., 2013; Kaiser et al., 2020; Guss et al., 2021; Baker et al., 2022; Beattie et al., 2016)—which provide only coarse, indirect signals of rollout quality. Genie (Bruce et al., 2024; Parker-Holder et al., 2024) decouples world model learning and agent training, but its evaluation still emphasizes visual quality and control, without probing semantic or causal fidelity. Cosmos (Agarwal et al., 2025) proposes a structured protocol that combines FID/FVD with structure-from-motion-based 3D consistency checks and human ratings on instruction following, object permanence, and visual verity. While insightful, this approach is tied to simulator-specific infrastructure and requires costly manual comparison. Human-in-the-loop protocols such as the Video Generation Arena (Analysis, 2024) also rely on pairwise comparison to assess rollout quality. These methods, though informative, are expensive and hard to scale.
@@
Evaluation Metrics and Protocols for Visual Generation. Early evaluations of generative models relied on full-reference metrics such as PSNR and SSIM (Wang et al., 2004), which capture pixel-level and perceptual similarity but are sensitive to spatial misalignments and fail to reflect semantic fidelity. To address this, distributional metrics like Inception Score (IS) (Salimans et al., 2016), Fréchet Inception Distance (FID) (Heusel et al., 2017), and Kernel Inception Distance (KID) (Binkowski et al., 2018). Other proposals such as PPL (Karras et al., 2019), Parzen likelihoods (Goodfellow et al., 2014), and HYPE (Zhou et al., 2019) attempt to quantify perceptual smoothness or human realism, but remain focused on static images. For video generation, FVD (Unterthiner et al., 2018) generalizes FID using I3D features (Carreira & Zisserman, 2017), introducing a motion-aware distributional baseline. Yet, FVD also lacks semantic grounding and does not account for causal structure or goal alignment. To improve semantic grounding, metrics based on text-image alignment have been proposed. CLIPScore (Hessel et al., 2021) and CLIPSIM (Wu et al., 2021) compute similarity between generated visuals and textual or visual references using CLIP embeddings (Radford et al., 2021), while Jayasumana et al. (2024) extend this to distributional comparisons via MMD. However, all operate at the frame level. Structured evaluation protocols using vision-language reasoning have also emerged. VQA Accuracy (Mañas et al., 2024) uses LLMs to score answers on static image questions, and VQAScore (Lin et al., 2024) probes alignment via templated binary queries. Lee et al. (2024) propose VLM evaluator to evaluate other VLMs responses given user criteria. These approaches introduce task structure but remain limited to single-frame evaluation. Recent text-to-video (T2V) benchmarks such as EvalCrafter (Liu et al., 2024b), VBench (Huang et al., 2024), and DEVIL (Liao et al., 2024) introduce curated prompts and metrics covering text alignment, motion realism, and perceptual quality. While these protocols push forward evaluation of open-ended video generation, they lack timestamp-level action grounding.
@@
Vision-Language Model Adaptation. VLMs have emerged as powerful tools for multimodal understanding, demonstrating strong performance across tasks such as captioning, retrieval, visual question answering, and instruction following (Hendriksen et al., 2022; Li et al., 2023; Driess et al., 2023; Chen et al., 2023; Wang et al., 2024; Abdin et al., 2024; Liu et al., 2024a; Deitke et al., 2024; McKinzie et al., 2024). Adaptation approaches can be broadly categorized into prompt-level and weight-level methods. One prominent prompt-level adaptation techniques is prompt tuning, which injects task information directly into the input space (Miyai et al., 2023; Zhou et al., 2024; Wu et al., 2024a), and in-context learning (ICL), where models such as GPT-3 (Brown et al., 2020) and Flamingo (Alayrac et al., 2022) condition on task demonstrations at inference time without updating parameters. Retrieval-augmented generation (RAG) (Lewis et al., 2020) combines parametric models with non-parametric memory, and multimodal variants incorporate external visual or auditory context (Hu et al., 2023b; Chen et al., 2022a). While lightweight, these approaches are limited in their ability to model temporal dependencies or align with structured rollouts. Weight-level adaptation enables stronger domain alignment but incurs higher computational cost. Full finetuning remains effective yet costly, while partial finetuning (Ye et al., 2023) offers a trade-off by updating only selected layers. Parameter-efficient finetuning (PEFT) provides a scalable alternative and can be grouped into low-rank and adapter-based strategies (Han et al., 2024). Low-rank methods, such as LoRA (Hu et al., 2022), inject rank-constrained updates into frozen layers. Recent extensions improve upon this via weight decomposition (Liu et al., ), quantization-aware adaptation (Dettmers et al., 2023; Xu et al., 2024), mixture-of-experts routing (Wu et al., 2024b), and long-context support (Chen et al., ). Adapter-based methods insert lightweight modules between frozen layers to enable modular adaptation with minimal overhead (Luo et al., 2023; Zhao et al., 2024). A parallel line of work investigates multimodal few-shot learning. Frozen (Tsimpoukelli et al., 2021) was among the first to explore this setting, followed by works combining prompting and ICL for improved sample efficiency (Jin et al., 2022; Song et al., 2022), and works introducing a learnable meta-mapper to bridge frozen VLM components for few-shot meta-learning (Najdenkoska et al., 2023).
Our Focus. While prior efforts have explored related challenges, none directly address the evaluation of the structured, action-conditioned fidelity and semantics of world model rollouts using adapted VLM. To this end, we introduce: (i) an evaluation protocol for world model rollouts, targeting fine-grained, temporally grounded assessment of semantic fidelity; (ii) universe, a VLM-based method to support the protocol. We validate its alignment with human judgments and demonstrate its scalability and semantic sensitivity across rollout conditions.
@@ 3 Methodology
Evaluation Protocol. We define two structured recognition tasks: (i) Action Recognition (AR): Assesses whether generated sequences accurately reflect the effects of agent actions given the segment; (ii) Character Recognition (CR): Evaluates whether entities maintain consistent identity and appearance across time. Each task is framed as a visual QA problem: the evaluator receives a sequence of frames and a natural language prompt (binary, multiple-choice, or open-ended), and generates a textual response. Outputs are scored using Exact Match (EM) and ROUGE-F1 (ROUGE), capturing both literal and semantic alignment with the reference answer. Metric details are in Appendix E.2.
Dataset Construction. Effective VLM adaptation for rollout evaluation requires a dataset that (i) captures realistic human behavior in interactive environments, and (ii) aligns with prior work in simulated settings to support comparability and reproducibility. To satisfy these constraints, we partnered with Ninja Theory and curated a dataset from both internal and public Bleeding Edge gameplay recordings, focusing on the Skygarden evnironment (Kanervisto et al., 2025). This dataset provides high visual and behavioral diversity (Pearce et al., 2025), includes a publicly available evaluation split, and is closely aligned with prior work in the domain (Kanervisto et al., 2025; Pearce et al., 2025; Tot et al., 2025; Sharma et al., 2024; Devlin et al., 2021), enabling fair comparison.
Data preparation proceeds in three stages: (i) Preprocessing:Segment gameplay into 14-frame clips with synchronized video, control logs, and metadata; (ii) Description Generation:Convert structured annotations (e.g., actions, agent states) into natural language summaries; (iii) Question-Answer Pair Construction:Generate six complimentary QA pairs per clip (binary, multiple-choice, and open-ended) spanning both AR and CR tasks. The final dataset contains 32.453 training clips and 8.113 validation clips, yielding 194.718 and 48.678 QA pairs, respectively. See Appendix D for details.
Model Architecture. We adapt a model from the PaliGemma family (Beyer et al., 2024; Steiner et al., 2024), consisting of a vision encoder $\mathcal{M}_{V}$, a projection head $\mathcal{M}_{P}$, and a language decoder $\mathcal{M}_{L}$. Based on initial zero-shot evaluations (Appendix G.2), we use a single configuration for all experiments—PaliGemma 2 3b, which includes a 2B-parameter Gemma 2 decoder pretrained on 2T tokens. Input frames are resized to $224\times 224$ and tokenized into 256 patches each. Model achitecture details are in Appendix E.1.
Each model input sequence $S=\{S_{\mathcal{I}},S_{\mathcal{T}}^{\text{PREF}},S_{\mathcal{T}}^{\text{SUFF}}\}$ consists of: visual tokens $S_{\mathcal{I}}$ from $k$ frames, a textual prefix $S_{\mathcal{T}}^{\text{PREF}}$ containing the task-language cue and question, and a suffix $S_{\mathcal{T}}^{\text{SUFF}}$ with the expected answer (used only during training). This format allows the decoder to attend jointly over visual and textual context. Full prompt details are provided in Appendix E.1.
@@
Fine-Tuning Configurations. We compare five adaptation strategies varying in parameter count and modularity: (i) Zero-shot prompting:No tuning; model is prompted directly. (ii) Full fine-tuning:All parameters $\theta=\theta_{V}\cup\theta_{P}\cup\theta_{L}$ are updated end-to-end. (iii) Dual-component fine-tuning:Two of three modules are trained (e.g., $\theta_{P}\cup\theta_{L}$). (iv) Single-component fine-tuning:Only one module—vision, projection, or language—is updated. (v) Parameter-efficient fine-tuning:We apply LoRA (Hu et al., 2022) adapters to attention and MLP layers in vision and language components: $\mathbf{W}\leftarrow\mathbf{W}+\frac{\alpha}{r}\mathbf{A}\mathbf{B},\quad\alpha=8,\;r\in\{8,16,32,48,64\}$.
@@
universe: UNIfied Vision-language Evaluator for Rollouts in Simulated Environments. We distill our empirical findings into universe, a lightweight and scalable adaptation method for temporally grounded evaluation of world model rollouts using VLMs. Designed for constrained compute and limited supervision, universe delivers strong generalization across our evaluation protocol using a single, partially tuned model. The method combines three main components: I Partial fine-tuning:We update only the projection head ($\theta_{P}$), training just 0.07% of model parameters. Despite this minimal footprint, it achieves the second-best performance among all strategies—trailing only vision encoder tuning, which requires $\sim$11% of parameters and incurs significantly higher compute cost. II Efficient frame sampling:Each input sequence includes $k=8$ frames sampled uniformly from a 14-frame rollout. This sparsity-aware strategy maintains long-range temporal structure while reducing token count and enabling efficient batching. III Mixed supervision:We train on a hierarchical mixture of tasks and QA formats. The task distribution favors Action Recognition ($\alpha_{\texttt{AR}}=0.8$) as it converges slower. Within each task, we emphasize open-ended questions ($\beta_{\texttt{OE}}=0.8$), while maintaining smaller proportions of binary ($\beta_{\texttt{binary}}=0.15$) and multiple-choice ($\beta_{\texttt{MC}}=0.05$) examples.